An investor bought $100,000 worth (roughly $750,000 face value!) of 30-year zero coupon Treasury STRIPS priced to yield 7%. One month later the yield on the STRIPS had risen to 8%. Use the modified duration approximation to estimate the approximate new market value shown when the investor looked at the next monthly statement.

  • A. $108,000

  • B. $99,000

  • C. $92,600

  • D. $72,000
Answer #1

Since Zero coupon bond; Macaulay duration =30 years

Modified duration =30/(1+7%) =28 years

This means when yield increase by 1% ; price decreases by 28.037%

New price= 100000*(1-.28) =72000